Can you explain to me the process on graphing freactions ex(-5/2,3/2) on a rectangular coordinate system.
Answer by checkley77(12844)

x=-5/2=-2.5 units on the x axis.
y=3/2=1.5 units on the y axis.
Plot these 2 points & then draw a line through them.